OpenStudy (anonymous):

[6.05] An airplane flew 3 hours with a 40 mph head wind. The return trip with a tail wind of the same speed took 2 hours. Find the speed of the plane in still air. 180 mph 200 mph 220 mph 240 mph

OpenStudy (anonymous):

is the answer 240?

OpenStudy (kropot72):

Let s = still air speed of plane Distance traveled on outward trip = (s - 40) * 3 = d Distance traveled on return trip = (s + 40) * 2 = d 3(s - 40) = 2(s + 40) 3s - 120 = 2s + 80 Can you solve for s ?

OpenStudy (anonymous):

yes and i got 40 as my answer

OpenStudy (kropot72):

3s - 120 = 2s + 80 Subtract 2s from each side (3s - 2s) -120 = (2s - 2s) + 80 s -120 = 0 + 80 Add 120 to each side s - (120 +120) = 80 + 120 s = 80 + 120 = ? mph

OpenStudy (anonymous):

200

OpenStudy (kropot72):

200 mph is the correct choice.

OpenStudy (anonymous):

thanks :)

OpenStudy (kropot72):

You're welcome :)